# Why Isn’t Relational Learning Taking Over the World?

### v4

@@ Introduction
The simplest form of relational data is exemplified by collaborative filtering datasets such as used in the Netflix challenge (Bell and Koren 2007) and Movielens datasets (Harper and Konstan 2015). Movielens contains rating datasets containing user-movie-rating-timestamp tuples, where the user and movie are identifiers represented by arbitrary numbers. There are also tables for user properties and movie properties, including links into the Internet Movie Database (IMDB). An even simpler version is a rated dataset that ignores the actual ratings, and is just a user-movie dataset (perhaps with timestamps). To predict rating (or rated) from just the rating dataset, you can use latent properties (embeddings) of users and of movies to predict ratings. Representing latent properties as a vector, matrix factorization (Koren, Bell, and Volinsky 2009) learns about movies and users, without learning general knowledge. Another possible technique is to do a self-join, which can result in a user-user table or a movie-movie table.
@@
  • Do unique identifiers denote different entities? This is known as the unique names assumption and features in logic programming and most databases. Identifying equality (two descriptions describing the same entity) arises in record linkage (Fellegi and Sunter 1969) and citation matching (Pasula et al. 2003). It is a big problem in medicine where drug dispensers need to decide whether someone asking for drugs is the same person who asked previously.
@@
(Subject, verb, object) triples are the basis for knowledge graphs where the subject is an entity, the verb is a relation or property and the object is an entity or a datatype value, such as a number or a string. A collection of triples can be treated as a labelled directed graph where the entities and datatype values are nodes. There is an arc from the subject to the object, labelled with the verb. Triples are universal representations of relations: any relation can be represented as subject-verb-object triples (See e.g., Poole and Mackworth 2023, Section 16.1).
One large public knowledge graph is Wikidata (Vrandečić and Krötzsch 2014). Figure 2 (a) shows some tables representing the football (soccer) player Christine Sinclair. She is the player (man or woman) with the most goals in international play. Figure 2 (b) shows the corresponding Wikidata knowledge graph. In Wikidata, Christine Sinclair is represented by the identifier Q262802. The country Canada is represented by the identifier Q16.
@@
member team start_time #matches #goals
@@
Figure 2: Facts about football player Christine Sinclair (Q262802): (a) in tabular form and (b) part of the corresponding Wikidata knowledge graph; from Figure 16.1 of Poole and Mackworth (2023)
@@
The second way to convert relations to triples is to use (row, column, value) triples. The row can be a primary key if there is one, or is a new identifier – the “reified” entity. Reify means “make into an entity” (or “thingify”). In Figure 2 (b), Q262802-3B1AE42E and Q262802-9c5c267f are reified entities. The number of facts about a reified entity is the number of columns in the table, which is typically just a handful. To convert to the original relation, one can do a database join of the triples on the reified entities and project onto the other properties (which is how I generated the table from the Wikidata triples). Reification does not lose information, but make the representation more modular: other properties, including meta-properties (such as the author) can be easily added to a knowledge graph.
@@
Note that graph learning (Hamilton 2020) and relational learning are isomorphic problems. A graph can be represented as a tail-head-label “arc” relation. Relations can be represented as knowledge graphs. These are complementary fields. It is not clear that a typical relational database is anything like a typical graph. Knowledge graphs tend to be much more heterogenous than graph benchmarks (Bechler-Speicher et al. 2025); the only general structure is that induced by reification. Many of the questions of interest are different, for example, there tends to be only a single knowledge graph so graph matching (Roy et al. 2025) is not really a problem of interest. Most of the issues in this paper do not apply to graph learning in general.
@@
One of the most used knowledge graphs in the literature is FB15k (Bordes et al. 2013b), a large subset of Freebase. Freebase (Bollacker et al. 2008) was a free and open knowledge base developed from 2007-2016. It was acquired by Google, who donated the contents to Wikidata, and use another branch as the Google knowledge base. In building FB15k, Bordes et al. (2013b) selected entities and relations that appeared in more than 100 tuples. They also created a larger knowledge graph, FB1M, with the most frequently occurring 1 million entities, which has been used much less in subsequent research.
@@
Figure 3: Some of the test triples in FB15k. The subject and object entities have their identifier replaced with the name. The verb (property) names have been abbreviated.
@@
We [^1] Bahare Fatemi, Mehran Kazemi, Ali Mehr and Martin Wang. See https://cs.ubc.ca/˜poole/res/FB15k-triples.txt for example triples for all 1344 predicates. created English version of the test set that, together with the plausible examples that existing methods ranked highly, could be used to crowdsource ground truth. The ground truth could then be used to evaluate prediction methods. However, we decided determining whether a tuple was true wasn’t a skill people could do quickly and accurately, even when they had access to Wikipedia, Google and other online sources. Positive examples were straightforward, but determining something is false requires guessing.
Toutanova and Chen (2015) noticed that a predictor could do reasonably well on FB15k by just determining that two predicates were (approximately) equivalent or inverses. To determine whether a predictor was doing more than this, they created FB15k-237 with the inverse functions removed (it would remove one first two triples of Figure 3). FB15k-237 isn’t better than FB15k; they test different things. Together they provide interesting evidence: the difference tells us whether a predictor can recognize inverses; the performance on FB15k-237 tells us whether the method can exploit other regularities. We would want a predictor that can determine inverses, but would like it to do more than that.
@@
Big data often implies more entities, but less data, on average, about each entity. As entities are added, they start off as stubs with few tuples, resulting in a long tail of entities about which little is known. As more detail is added about an entity, there are more associated reified entities with inherently few tuples (see Appendix A).
Another standard dataset is WN18 (Bordes et al. 2013a), based on WordNet (Miller 1995), a large lexical database of English. The entities are words, and the relations include hyponym (subclass), hypernym (superclass), part_of and type, with loosely connected hierarchies for nouns, verbs, adjectives and adverbs. Most of the relations are transitive. I would not expect that words are prototypical entities.
@@
In relational datasets with no negative statements, the prediction that optimizes log loss is to predict everything is true, with a loss of zero on any data set that only includes positive statements. However for a test set that includes statements that should be false, this prediction has an infinite test loss. In training, the lack of negative statements is generally handled by a form of contrastive learning: adding random tuples and using these as negative examples (the few that might be correct are noise). The number of negative versus positive examples provides input that is just fiction. Unlike a prior probability, this input is not overwhelmed by observed data.
@@
  • parameter sharing or weight tying; namely “logical variables” and “quantification” in logic approaches (Poole 1993; De Raedt, Kimmig, and Toivonen 2007; Domingos and Lowd 2009), exchangeability [^2] Identifiers can be exchanged and the probability distribution should not change. Exchangeability of identifiers does not imply exchangeability of random variables (de Finetti 1974). (Niepert and Van den Broeck 2014), or the “convolutional” in convolutional graph neural networks (Schlichtkrull et al. 2018).
@@
The most common way to evaluate a knowledge graph predictor is to use ranking: given test triple $(s,v,o)$, ask the predictor to rank the entities related to a given subject $s$ and verb $v$, i.e., to predict which triples of the form $(s,v,?)$ are true (and similarly to rank the subjects for a given verb and object). Given a total ordering of predictions $o_{1},o_{2},o_{3},\dots$, the rank is the $i$ such that $o=o_{i}$. In filtered prediction (Bordes et al. 2013b), the $o_{i}$ that are known to be true (because the corresponding triple is in the training or test set) are removed from the ordering.
The mean rank is one possible measure but is dominated by a few very large ranks, and is not used much any more. A common measure is “hit-at-kk” for some integer $k$ (such as 1, 3 or 10), the proportion of test cases where the rank of the correct answer is $k$ or less. The $k$ in hit-at-$k$ seems like an arbitrary parameter. A method that is common and seems to not have arbitrary parameters is the mean reciprocal rank (MRR) (Radev et al. 2002). This weights the $k$th rank as $1/k$. It can be interpreted as an agent who always considers the first item, considers the second item half the time, the 3rd item 1/3 of the time on so on. However this interpretation is problematic (see Appendix B).
@@
  • Some questions cannot be asked such as “who is the Pope married to?” (where the answer is “nobody”).
  • Asking $(s,v,?)$ leaks information about the test set. Asking who $e$ is married to tells us that $e$ is a person, an adult and not the Pope. Knowing that someone rated a movie, as can be obtained from asking for a prediction for the rating of a movie, is valuable information that can be exploited to produce better predictions. One could imagine a task where the test set is meant to provide information about an entity, as a prompt to an LLM might, and asks questions about that entity. However for many tasks, questions should not provide information. Whether leaking information about the test set is appropriate depends on the downstream application.
  • Some queries are trivial; for example, with fewer than 10 football (soccer) positions and most teams having all of them, asking for a position in a team (see Figure 3), any list of all positions will give a hit-at-10.
  • Some queries are almost impossible, for example asking which soccer team has someone playing “forward” – one of the queries of the FB15k test set – means listing all soccer teams. Guessing which one the query is about is almost impossible. You would expect that an omniscient agent – one who knows everything – would go well in any reasonable evaluation. However, an omniscient agent who knows all soccer teams and all players and their positions, would not be able to guess what team the query was about.
@@
For incomplete knowledge bases, we need to explicitly consider why data is missing. Having test examples that are a random subset of all of the tuples, does not provide a surrogate to real-world problems where data is invariably not missing at random (Rubin 1976). I doubt there are any Taylor Swift officially released albums missing from Wikidata, but for most of the recording artists on Wikidata most of their recordings are missing. Probabilistic models of missing data (Marlin et al. 2011; Mohan, Pearl, and Tian 2013) tend to be heavyweight, requiring detailed modelling of why data is missing. This is difficult when virtually all of the tuples are missing. We made an attempt to build a lightweight model of missing data (Poole, Mehr, and Wang 2020) with the aim of extended it to relational models.
@@
Aggregation (e.g., predicting gender of a user from the movies they watched) is the Achilles heel of relational learning. Most models either have a built-in aggregation such as noisy-or for probabilistic logic programs (Poole 1993) or logistic regression for weighted logical formula models (Domingos and Lowd 2009; Kazemi et al. 2014), use a distribution of related entities in relational attention (Velickovic et al. 2018; Wang et al. 2019) or use explicit operations such a sum or mean (Schlichtkrull et al. 2018). They are problematic theoretically (e.g., asymptoticly) (Poole et al. 2014; Buchman and Poole 2015; Adam-Day et al. 2024) and in practice (Kazemi et al. 2017). They need to work with no related entities and as the number of related entities goes to infinity. The models either implicitly assume the related entities provide independent evidence (as do sum, noisy-or and logistic regression), or act the same as if there were one or few related entities (as do max, average and attention). Determining whether the evidence is independent (so can be accumulated) or is dependent (in which case extra evidence can be discounted) is difficult, and invariably ignored.
@@ Acknowledgements
Appendix A Appendix A: Proportion of Wikidata entities with few tuples
The aim of this appendix is to estimate the proportion of entities in Wikidata that are the subject of few tuples. There is very long tail of entities about which there are few tuples, and presumably are stubs that will be expanded in the future, so it may be unfair to count those. Instead, I approximate the number of reified entities per top-level entity. I could not download all of Wikidata, so as a sample I selected the entities of the Wikipedia featured articles of March 2025 (https://en.wikipedia.org/wiki/Wikipedia:Today%27s˙featured˙article/March˙2025) that had Wikidata pages. The entities were diverse including Hughie Ferguson, a Scottish footballer in 1910s and 20s, Pulgasari, a North Korean film from 1985, the contemporary ballet Flight Pattern, Steele’s Greenville expedition in the American Civil War, the sun, and a fictional character in the medical drama series House.
For each entity (including reified entities), Wikidata allows users to download triples about that entity. For reified entities, the triples downloaded are the triples of the subject of the reification. Define the local entities for entity $e$ to be those that appear as a subject in a triple for the page for $e$ and for which Wikidata downloads the page for $e$. For each of the selected entities, I counted the number of local entities that are the subject of fewer than 10 triples and the number of entities that are the subject of 10 or more triples. See Figure 4. The number ranged from 5 for History of infant schools in Great Britain and 6 for Your Girl (a 2005 song by Mariah Carey), to 841 for the Sun and 520 for the composer and conductor Pierre Boulez. The entities with the lowest number of tuples are just stubs; note that for Q133116210, the 11 tuples included the name and labels in both English and Portuguese. The average number of local entities was 132 (or 108 if the top and bottom outliers were removed).
Wikidata ID # tuples <10<10 $\geq 10$ name
@@ Appendix A Appendix A: Proportion of Wikidata entities with few tuples
Figure 4: For each top-level entity: the Wikidata ID, the number of tuples with the entity as subject, the number of local entities that are the subject of fewer than 10 tuples, the number with 10 or more tuples, and the English name of the entity. Sorted by number of tuples. These numbers were from 2025-08-19; they change over time as the knowledge base evolves. The Python code to generate the contents of this table is available from the author.
The entities that are the subject of 10 or more tuples (other than the top-level entity) are refied entities from relations with a large number of arguments; the maximum of which is 17. About 1.5% of the entities are the subject of 10 or more triples. Fewer than 0.75% of the entities are the subject of 18 or more triples.
As the number of triples increases for an entity, so does the number of entities with few triples. Thus, as a knowledge graph gets bigger and more filled out, the proportion of entities with few triples increases. This goes against the idea that we just need more data. There are also top-level entities with few triples.
Note that Wikidata did not include entities for all of the topics of the Wikipedia features articles. The March 12 wikipedia article was for the 2020 season for Seattle Sounders FC, which was not an entity in Wikidata, so we used Seattle Sounders FC (Q632511). There was no Wikidata entry for the March 14 page on the five-pound British gold coin, or the March 31 entry for Apollo 15 postal covers incident. These were omitted. There were 2 entities for “geography of Ireland”, one for the country and one for the island; the featured article was for the island, so that is the entity used.

### v5

@@ Introduction
The simplest form of relational data is exemplified by collaborative filtering datasets such as used in the Netflix challenge (Bell and Koren 2007) and Movielens datasets (Harper and Konstan 2015). Movielens contains rating datasets containing user-movie-rating-timestamp triples, where the user and movie are identifiers represented by arbitrary numbers. There are also tables for user properties and movie properties, including links into the Internet Movie Database (IMDB). An even simpler version is a rated dataset that ignores the actual ratings, and is just a user-movie dataset (perhaps with timestamps). To predict rating (or rated) from just the rating dataset, you can use latent properties (embeddings) of users and of movies to predict ratings. Representing latent properties as a vector, matrix factorization (Koren, Bell, and Volinsky 2009) learns about movies and users, without learning general knowledge. Another possible technique is to do a self-join, which can result in a user-user table or a movie-movie table.
@@
  • Do different identifiers denote different entities? This is known as the unique names assumption and features in logic programming and most databases. Identifying equality (two descriptions describing the same entity) arises in record linkage (Fellegi and Sunter 1969) and citation matching (Pasula et al. 2003). It is a big problem in medicine where drug dispensers need to decide whether someone asking for drugs is the same person who asked previously.
@@
(Subject, verb, object) triples are the basis for knowledge graphs. The subject is an entity. The verb is a relation or property. The object is an entity or a datatype value, such as a number or a string. A collection of triples can be treated as a labelled directed graph where the entities and datatype values are nodes. There is an arc from the subject to the object, labelled with the verb. Triples are universal representations of relations: any relation can be represented as subject-verb-object triples (See e.g., Poole and Mackworth 2023, Section 16.1).
One large public knowledge graph is Wikidata (Vrandečić and Krötzsch 2014). Table 1 shows tables about the football (soccer) player Christine Sinclair. She is the player (man or woman) with the most goals in international play. Figure 2 shows part of the corresponding Wikidata knowledge graph. In Wikidata, Christine Sinclair is represented by the identifier Q262802. The country Canada is represented by the identifier Q16.
@@
member team start #matches #goals
@@
Table 1: Facts about football player Christine Sinclair (Q262802) in tabular form
Figure 2: Part of the Wikidata knowledge graph about Christine Sinclair; from Figure 16.1 of Poole and Mackworth (2023)
@@
The second way to convert relations to triples is to use (row, column, value) triples. The row can be a primary key if there is one, or is a new identifier – a reified entity. Reify means “make into an entity” (or “thingify”). In Figure 2, Q262802-3B1 and Q262802-9c5 are reified entities. The number of facts about a reified entity is the number of columns in the table, which is typically just a handful. To convert to the original relation, one can do a database join of the triples on the reified entities and project onto the other properties (which is how I generated the tables from the Wikidata triples). Reification does not lose information. It makes the representation more modular: other properties, including meta-properties (such as the author) can be easily added to a knowledge graph.
@@
Note that graph learning (Hamilton 2020) and relational learning are isomorphic problems. A graph can be represented as a tail-head-label arc relation. Relations can be represented as knowledge graphs. These are complementary fields. It is not clear that a typical relational database is anything like a typical graph. Knowledge graphs tend to be much more heterogenous than graph benchmarks (Bechler-Speicher et al. 2025); the only general structure is that induced by reification. Many of the questions of interest are different. For example, there tends to be only a single knowledge graph so graph matching (Roy et al. 2025) is not really a problem of interest. Most of the issues in this paper do not apply to graph learning in general.
@@
One of the most used knowledge graphs in the literature is FB15k (Bordes et al. 2013), a large subset of Freebase. Freebase (Bollacker et al. 2008) was a free and open knowledge base developed from 2007-2016. It was acquired by Google, who donated the contents to Wikidata, and use another branch as the Google knowledge base. In building FB15k, Bordes et al. (2013) selected entities and relations that appeared in more than 100 triples. They also created a larger knowledge graph, FB1M, with the most frequently occurring 1 million entities, which has been used much less in subsequent research.
@@
Figure 3: Some of the triples in FB15k. The subject and object entities have their identifier replaced with the name. The verb (property) names have been abbreviated.
@@
We [^1] Bahare Fatemi, Mehran Kazemi, Ali Mehr and Martin Wang. See https://cs.ubc.ca/˜poole/res/FB15k-triples.txt for example triples for all 1344 predicates. created English version of the test set that, together with the plausible examples that existing methods ranked highly, could be used to crowdsource ground truth. The ground truth could then be used to evaluate prediction methods. However, we decided determining whether a triple was true wasn’t a skill people could do quickly and accurately, even when they had access to Wikipedia, Google and other online sources. Positive examples were straightforward, but determining something is false requires guessing.
Toutanova and Chen (2015) noticed that a predictor could do reasonably well on FB15k by just determining that two predicates were (approximately) equivalent or inverses. To determine whether a predictor was doing more than this, they created FB15k-237 with the inverse functions removed (it would remove one of the first two triples of Figure 3). FB15k-237 isn’t better than FB15k; they test different things. Together they provide interesting evidence: the difference tells us whether a predictor can recognize inverses; the performance on FB15k-237 tells us whether the method can exploit other regularities. We would want a predictor that can determine inverses, but would like it to do more than that.
@@
Big data often implies more entities, but less data, on average, about each entity. As entities are added, they start off as stubs with few triples, resulting in a long tail of entities about which little is known. As more detail is added about an entity, there are more associated reified entities with inherently few triples (see Appendix A).
Another standard dataset is WN18 (Bordes et al. 2013), based on WordNet (Miller 1995), a large lexical database of English. The entities are words, and the relations include hyponym (subclass), hypernym (superclass), part_of and type, with loosely connected hierarchies for nouns, verbs, adjectives and adverbs. Most of the relations are transitive. I would not expect that words are prototypical entities.
@@
In relational datasets with no negative statements, the prediction that optimizes log loss is to predict everything is true, with a loss of zero on any data set that only includes positive statements. However for a test set that includes statements that should be false, this prediction has an infinite test loss. In training, the lack of negative statements is generally handled by a form of contrastive learning: adding random triples and using these as negative examples (the few that might be correct are noise). The number of negative versus positive examples provides input that is just fiction. Unlike a prior probability, this input is not overwhelmed by observed data.
@@
  • parameter sharing or weight tying; namely logical variables and quantification in logic approaches (Poole 1993; De Raedt, Kimmig, and Toivonen 2007; Domingos and Lowd 2009), exchangeability [^2] Identifiers can be exchanged and the probability distribution should not change. Exchangeability of identifiers does not imply exchangeability of random variables (de Finetti 1974). (Niepert and Van den Broeck 2014), or the convolutional in convolutional graph neural networks (Schlichtkrull et al. 2018).
@@
The most common way to evaluate a knowledge graph predictor is to use ranking: given test triple $(s,v,o)$, ask the predictor to rank the entities related to a given subject $s$ and verb $v$, i.e., to predict which triples of the form $(s,v,?)$ are true (and similarly to rank the subjects for a given verb and object). Given a total ordering of predictions $o_{1},o_{2},o_{3},\dots$, the rank is the $i$ such that $o=o_{i}$. In filtered prediction (Bordes et al. 2013), the $o_{i}$ that are known to be true (because the corresponding triple is in the training or test set) are removed from the ordering.
The mean rank is one possible measure but is dominated by a few very large ranks, and is not used much any more. A common measure is hit-at-$k$ for some integer $k$ (such as 1, 3 or 10), the proportion of test cases where the rank of the correct answer is $k$ or less. The $k$ in hit-at-$k$ seems like an arbitrary parameter. A method that is common and seems to not have arbitrary parameters is the mean reciprocal rank (MRR) (Radev et al. 2002). This weights the $k$th rank as $1/k$. It can be interpreted as an agent who always considers the first item, considers the second item half the time, the 3rd item 1/3 of the time on so on. However this interpretation is problematic (see Appendix B).
@@
  • Some questions cannot be asked, such as “who is the Pope married to?” (where the answer is “nobody”).
  • Asking $(s,v,?)$ leaks information about the test set. Asking who $e$ is married to tells us that $e$ is a person, an adult and not the Pope. Knowing that someone rated a movie, as can be obtained from asking for a prediction for the rating of a movie, is valuable information that can be exploited to produce better predictions. One could imagine a task that provides information about an entity, as a prompt to an LLM might, and asks questions about that entity. For many tasks, questions should not provide information. Whether leaking information about the test set is appropriate depends on the downstream application.
  • Some queries are trivial; for example, with fewer than 10 football (soccer) positions in FB15k and most teams having all of them, asking for a position in a team (see Figure 3), any list of all positions will give a hit-at-10.
  • Some queries are almost impossible, for example asking which soccer team has someone playing forward – one of the queries of the FB15k test set – means listing all soccer teams. Guessing which one the query is about is almost impossible. You would expect that an omniscient agent – one who knows everything – would go well in any reasonable evaluation. However, an omniscient agent who knows all soccer teams and all players and their positions, would not be able to guess what team the query was about.
@@
For incomplete knowledge bases, we need to explicitly consider why data is missing. Having test examples that are a random subset of all of the triples, does not provide a surrogate to real-world problems where data is invariably not missing at random (Rubin 1976). I doubt there are any Taylor Swift officially released albums missing from Wikidata, but for most of the recording artists on Wikidata most of their recordings are missing. Probabilistic models of missing data (Marlin et al. 2011; Mohan, Pearl, and Tian 2013) tend to be heavyweight, requiring detailed modelling of why data is missing. This is difficult when virtually all of the triples are missing. We made an attempt to build a lightweight model of missing data (Poole, Mehr, and Wang 2020) with the aim of extended it to relational models.
@@
Aggregation (e.g., predicting the gender of a user from the movies they watched) is the Achilles heel of relational learning. Most models either have a built-in aggregation such as noisy-or for probabilistic logic programs (Poole 1993) or logistic regression for weighted logical formula models (Domingos and Lowd 2009; Kazemi et al. 2014), use a distribution of related entities in relational attention (Velickovic et al. 2018; Wang et al. 2019) or use explicit operations such a sum or mean (Schlichtkrull et al. 2018). They are problematic theoretically (e.g., asymptoticly) (Poole et al. 2014; Buchman and Poole 2015; Adam-Day et al. 2024) and in practice (Kazemi et al. 2017). They need to work with no related entities and as the number of related entities goes to infinity. The models either implicitly assume the related entities provide independent evidence (as do sum, noisy-or and logistic regression), or act the same as if there were one or few related entities (as do max, average and attention). Determining whether the evidence is independent (so can be accumulated) or is dependent (in which case extra evidence can be discounted) is difficult, and invariably ignored.
@@ Acknowledgements
Appendix A Appendix A: Proportion of Wikidata entities with few triples
The aim of this appendix is to estimate the proportion of entities in Wikidata that are the subject of few triples. I could not download all of Wikidata, and it is not clear what a random sample is. Picking entities at random results in nearly all of them being in very long tail of entities with very few triples, but this may not be representative of the triples. Wikidata does not provide a tool to access a random triple. To get a representative (if not random) sample, I selected the entities of the Wikipedia featured articles of March 2025 (https://en.wikipedia.org/wiki/Wikipedia:Today%27s˙featured˙article/March˙2025) that had Wikidata pages. The entities were diverse including Hughie Ferguson, a Scottish footballer in 1910s and 20s, Pulgasari, a North Korean film from 1985, the contemporary ballet Flight Pattern, Steele’s Greenville expedition in the American Civil War, the sun, and a fictional character in the medical drama series House. I tried a few other collections of entities and the results were similar. This is not meant to be a statistically valid study.
For each entity (including reified entities), Wikidata allows users to download triples about that entity. For reified entities, the triples downloaded are the triples of the subject of the reification. Define the local entities for entity $e$ to be those that appear as a subject in a triple for the page for $e$ and for which Wikidata downloads the page for $e$. For selected entity, I counted the number of local entities that are the subject of fewer than 10 triples and the number of entities that are the subject of 10 or more triples. See Figure 4. The number ranged from 5 for History of infant schools in Great Britain and 6 for Your Girl (a 2005 song by Mariah Carey), to 520 for the composer and conductor Pierre Boulez, and 841 for the Sun. The entities with the lowest number of triples are just stubs; note that for Q133116210, the 11 triples included the name and labels in both English and Portuguese. The average number of local entities was 134 (or 108 if the top and bottom outliers were removed).
Wikidata ID # triples <10<10 $\geq 10$ name
@@ Appendix A Appendix A: Proportion of Wikidata entities with few triples
Figure 4: For each top-level entity: the Wikidata ID, the number of triples with the entity as subject, the number of local entities that are the subject of fewer than 10 triples, the number with 10 or more triples, and the (truncated) English name of the entity. Sorted by number of triples. These numbers were from 2025-08-19; they change over time as the knowledge base evolves. The Python code to generate the contents of this table is available from the author.
The entities that are the subject of 10 or more triples (other than the top-level entity) are refied entities from relations with a large number of arguments; the maximum of which is 17. About 1.5% of the entities are the subject of 10 or more triples. Fewer than 0.75% of the entities are the subject of 18 or more triples.
As the number of triples for an entity increase, so does the number of associated (local) entities with few triples. As a knowledge graph gets bigger and more filled out, the proportion of entities with few triples increases. This goes against the idea that we just need more data.
Wikidata did not include entities for all of the topics of the Wikipedia features articles. The March 12 Wikipedia article was for the 2020 season for Seattle Sounders FC, which was not an entity in Wikidata, so we used Seattle Sounders FC (Q632511). There was no Wikidata entry for the March 14 page on the five-pound British gold coin, or the March 31 entry for Apollo 15 postal covers incident. These were omitted. There were 2 entities for “geography of Ireland”, one for the country and one for the island; the featured article was for the island, so that is the entity used.